# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v2

@@ Abstract
This paper examines the impact of racial discrimination in hiring on employment, wages, and wealth disparities between black and white workers. Using a labor search-and-matching model with racially prejudiced and non-prejudiced firms, we show that labor market frictions sustain discriminatory practices as an equilibrium outcome. These practices account for 44% to 52% of the average wage gap and 16% of the median wealth gap. Discriminatory hiring also amplifies unemployment and wage volatility for black workers, increasing their labor market risks over business cycles. Eliminating prejudiced firms reduces these disparities and improves the welfare of black workers, though it slightly decreases overall economic welfare.
@@
Keywords: search-and-matching, heterogeneous agents, racial inequality, unemployment, wealth distribution, business cycles
@@ 1 Introduction
Discrimination is often considered inefficient or a response to information asymmetry. [^1] See Becker (1957) and Lang and Spitzer (2020) for discussions on taste-based and statistical discrimination theories. Market competition and better information are expected to eliminate discrimination over time. Nonetheless, substantial empirical evidence shows that discriminatory practices persist across markets, perpetuating significant racial disparities. [^2] For example, Couch and Fairlie (2010); Biddle and Hamermesh (2013); Kuhn et al. (2020); Derenoncourt and Montialoux (2021); Lippens et al. (2023). Resume studies consistently reveal racial discrimination even when controlling for candidate qualifications (e.g., Bertrand and Mullainathan, 2004; Bertrand and Duflo, 2017; Kline et al., 2022). Gaps in labor income and wealth between black and white households have persisted long after the Civil Rights Movement (e.g., Cajner et al., 2017; Derenoncourt et al., 2023). Small and Pager (2020) argues that discrimination reinforces itself across domains and constitutes a form of market failure. Only recently has macroeconomic research begun exploring racial disparities. [^3] See, for example, Nakajima (2021), Aliprantis et al. (2023), Boerma and Karabarbounis (2021), Ganong et al. (2020), Lee et al. (2021). Yet, the role of discriminatory practices in shaping these disparities remains under-explored. Developing a general equilibrium framework to explain the persistence of discrimination and its consequences is crucial for understanding the macroeconomic impacts of racial disparities and informing effective policymaking.
@@
This paper constructs a labor search-and-matching framework to capture the persistence of hiring discrimination based on race. We show that hiring discrimination accounts for 44% to 52% of the black-to-white average wage gap and 16% of the racial median wealth gap. Additionally, hiring discrimination increases both unemployment and wage volatility among black workers. It creates a countercyclical racial unemployment gap and diverges cyclical wage patterns between races. The model incorporates two types of firms: prejudiced firms, which hire only white workers, and non-prejudiced firms, which hire all workers equally. We include additional racial differences in the model as differences in bargaining power and the likelihood of experiencing wealth destruction. While all workers compete for jobs, white workers have additional opportunities with prejudiced firms. Matched worker-firm pairs negotiate wages to maximize their joint surplus. Upon calibration, the model endogenously produces lower job-finding rates, higher unemployment rates, and lower bargained wages for black workers. Even without further financial frictions, the model generates significant racial disparities in wealth accumulation.
The main message of this paper is that labor market friction can sustain hiring discrimination as an equilibrium outcome. Our calibration shows that prejudiced firms incur higher vacancy posting costs to discriminate against black workers. As we gradually increase these costs, prejudiced firms slowly reduce their vacancy postings and employment, while their profits adjust to satisfy the free-entry condition. Ultimately, we more than double the benchmark calibrated posting cost until the entire expected value of entering as a discriminatory firm is eliminated. This indicates that prejudiced firms maintain sufficient profit to absorb higher vacancy posting costs, even under competitive market conditions.
@@
Moreover, hiring discrimination perpetuates the equilibrium economy through firms’ vacancy postings. Prejudiced firms compete with non-prejudiced firms for unemployed white workers, driving up wage rates for white workers across the economy. This wage pressure reduces the expected profits of non-prejudiced firms, leading them to post fewer vacancies overall. The resulting decline in non-prejudiced job openings directly limits employment opportunities for black workers, contributing to lower wages, higher unemployment rates, and adverse wealth accumulation for them.
In a counterfactual economy without prejudiced firms, the racial unemployment gap disappears, and the wage gap decreases significantly. Such a reduction in labor market disparity leads to asymmetric welfare changes. Black workers’ employment opportunities and wages increase because non-prejudiced firms expand their vacancy postings to fill positions left by prejudiced firms. As a result, black workers experience a welfare increase, with the most significant gains among high-productivity black workers. White workers, however, lose the exclusive job opportunities that prejudiced firms previously provided. Their average wage declines, and their unemployment rate rises. This leads to a moderate welfare decrease, especially among the lowest-productivity and lowest-wealth white workers. Due to asymmetrical job destruction shocks and vacancy posting costs, non-prejudiced firms cannot fully replace prejudiced firms. However, if we maintain the race-specific job destruction shocks, eliminating prejudiced hiring has a much smaller adverse impact on white workers.
We further introduce aggregate shocks to the economy and show that hiring discrimination generates distinct business cycle dynamics for black and white workers. Our model successfully replicates the observed countercyclical and volatile racial unemployment rate gap, as well as the more procyclical and volatile average wage rate for black workers. Upon removing prejudiced firms, the business cycle patterns of the racial unemployment rate gap disappear. The volatility and cyclicality of average wages for black and white workers become much more similar. This provides theoretical evidence to Cajner et al. (2017) that firm-side discrimination contributes to the business cycle differences between black and white workers, and non-market demographic factors alone do not fully account.
Lastly, we examine the associated business cycle welfare change between black and white workers when discriminatory hiring contributes to more volatile labor market outcomes for black workers. Not all workers benefit from a less volatile economy without prejudiced firms. Black workers gain an average of 0.02% in welfare, whereas white workers face a welfare loss of 0.01%. Different from the steady-state welfare values, among black workers, those with high productivity and high wealth gain the most. Among white workers, those with high productivity and high wealth lose the most, experiencing the largest welfare decreases within their group.
@@
This paper contributes to the ongoing discussion on the persistence of racial inequality by focusing on the disparate labor market conditions of white and black workers. Numerous studies have documented racial discrimination in pay and employment opportunities that persist despite policy interventions (e.g., Coate and Loury, 1993; Black, 1995; Rosén, 1997; Bertrand and Mullainathan, 2004; Manduca, 2018). Recent empirical research stresses that racial wage and employment gaps persist across education levels, skill sets, and cohorts. These disparities are not solely attributable to differences in individual qualifications (e.g., Pena, 2018; Cheng et al., 2019). Building on this insight, our paper develops a model in which discriminatory hiring practices endogenously perpetuate labor market inequalities. Our work adds to the literature by providing a theoretical framework that explains how hiring discrimination can be sustained as an equilibrium outcome in the labor market.
@@
While much of the literature focuses on static disparities, fewer studies document how macroeconomic fluctuations impact black and white workers differently. Notably, Couch and Fairlie (2010) demonstrate that black workers are disproportionately affected during recessions, being the last hired in upturns and the first fired in downturns. Biddle and Hamermesh (2013) find that the discriminatory wage gap is procyclical, widening during economic expansions. Similarly, Cajner et al. (2017) show that black workers experience higher unemployment rate volatility and are more likely to be involuntarily employed part-time. Daly et al. (2020) suggest that limited employment opportunities for black workers contribute to the growing racial earnings gap. Our paper adds to this body of work by evaluating the impact of hiring discrimination on the cyclical aspects of racial disparities, examining how discriminatory practices exacerbate volatility and cyclicality in employment and wages for black workers.
@@
An emerging strand of literature explores racial disparities in wealth accumulation. Studies such as Derenoncourt et al. (2023), Derenoncourt et al. (2022), Kuhn et al. (2020), Barsky et al. (2002), and McIntosh et al. (2020) document significant wealth gaps between black and white households. For example, Derenoncourt et al. (2023) provides a historical account of wealth segregation over the past 150 years, showing persistent and substantial disparities. Boerma and Karabarbounis (2021) and Aliprantis et al. (2023) analyze the impact of historical discrimination on earnings, bequests, and capital returns within steady-state models that do not consider aggregate risks. Furthermore, Ganong et al. (2020) demonstrates that income risks affect individuals differently across racial groups due to wealth differences. Building on these findings, Bartscher et al. (2021) and Lee et al. (2021) discuss how monetary policy can have disparate effects on workers of different racial backgrounds. Our paper contributes to this literature by showing that employer discrimination not only affects immediate labor market outcomes but also spills over into long-term wealth accumulation disparities between black and white workers.
@@
Germane to our project, Nakajima (2021) develops a search-and-matching model to examine how monetary policies exacerbate racial differences in the labor market. Our model explicitly distinguishes between prejudiced and non-prejudiced firms in the hiring process and directly examines how hiring discrimination impacts black workers and the broader economy. We analyze the spillover effects of discriminatory hiring practices on both labor and wealth disparities, offering a comprehensive theoretical understanding of these dynamics over the long run and across business cycles.
@@
The rest of the paper proceeds as follows. Section 2 lays out the theoretical framework. Section 3 discusses the calibration strategy. Section 4 examines the steady-state implications of racial discrimination. Section 5 provides the business cycle implications. Section 6 concludes the paper.
@@ 2.1 Environment
The model has a unit measure of workers, who are either black or white, $R\in\{bl,wh\}$. Two types of firms post vacancies to hire unemployed workers, subject to search and matching frictions. Prejudiced firms ($p$) actively discriminate against black workers and hire only white workers. Non-prejudiced firms ($np$) hire both black and white workers without discrimination. Prejudiced and non-prejudiced firms also experience type-specific job destruction shocks. Each firm posts one vacancy or hires one worker. Thus, firms and jobs are synonymous in our model. To ease exposition, we drop the time subscripts and use a prime symbol (’) to denote the variables in the next period.
Workers are either employed (with $p$ or $np$ firm) or unemployed. Those who become unemployed in the current period receive unemployment benefits and continue to receive them in the future with probability $P_{e}$. The workers face idiosyncratic productivity shocks $s$, following an AR(1) process $\log(s^{\prime})=\rho_{s}\log(s)+\epsilon_{s}$, with $\epsilon_{s}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{s}^{2})$. Workers also differ in their asset holdings. They can save using these risk-free assets to partially insure themselves against the labor market risks. Following Mukoyama (2013), workers also receive race-specific extreme wealth shocks with probability $\epsilon_{R}\in\{\epsilon_{bl},\epsilon_{wh}\}$. Upon realization of the shock, the worker loses all their wealth. Altogether, workers are heterogeneous across race ($R$), labor market status ($e$), idiosyncratic productivity ($s$), and wealth ($a$). The endogenous distribution of workers is summarized by $\mu(e,R,s,a)$. All endogenous aggregate variables depend on $\mu$. Individual workers may move around the distribution between periods; however, the overall distribution remains stationary. We omit the record of $\mu$ in the steady state and will reintroduce $\mu$ when augmenting the model with aggregate shocks in Section 5.
@@ 2.2 Labor market search and matching
The total number of unemployed workers, denoted as $u$, is the sum of unemployed black workers $(u_{bl})$ and unemployed white workers $(u_{wh})$. There are $v_{np}$ vacancies available in the non-prejudiced market, while the number of vacancies in the prejudiced market is $v_{p}$. Non-prejudiced firms search among both black and white unemployed workers, resulting in a non-prejudiced market tightness defined as $\theta_{np}=\frac{v_{np}}{u}$. In contrast, prejudiced firms only hire unemployed white workers, which gives the prejudiced market tightness $\theta_{p}=\frac{v_{p}}{u_{wh}}$. Following the works of Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018), unemployed workers and vacant firms match through a constant returns to scale matching function
$M(u,v)=\frac{uv}{(u^{\iota}+v^{\iota})^{1/\iota}},$ (1)
with $\iota>0$. As documented by Den Haan et al. (2000), this functional form ensures that the matching probabilities lie within 0 and 1. The probability for an unemployed worker to match with a vacant $np$ firm is $f(\theta_{np})=M(u,v_{np})/u=(1+\theta_{np}^{-\iota})^{-1/\iota}$, while the probability that a white unemployed worker matches with a vacant $p$ firm is $f(\theta_{p})=M(u_{wh},v_{p})/u_{wh}=(1+\theta_{p}^{-\iota})^{-1/\iota}$. Correspondingly, the probability of filling a vacant $np$ firm is $q(\theta_{np})=M(u,v_{np})/v_{np}=(1+\theta_{np}^{\iota})^{-1/\iota}$, while the probability of filling a vacant $p$ firm is $q(\theta_{p})=M(u_{wh},v_{p})/v_{p}=(1+\theta_{p}^{\iota})^{-1/\iota}$. In addition, non-prejudiced matches get separated with probability $\lambda_{np}$, while the prejudiced matches separate at the rate $\lambda_{p}$.

### v3

@@ Abstract
This paper examines the impact of racial discrimination in hiring on employment, wages, and wealth disparities between black and white workers. Using a labor search-and-matching model with racially prejudiced and non-prejudiced firms, we show that labor market frictions sustain discriminatory practices as an equilibrium outcome. These practices account for 57% of the racial unemployment gap, 48% of the average wage gap, and 16% of the median wealth gap. Discriminatory hiring also increases unemployment and wage volatility for black workers, increasing their labor market risks over the business cycle. Eliminating prejudiced firms reduces these disparities and improves the welfare of black workers as well as the overall economic welfare.
@@
Keywords: search-and-matching, heterogeneous agents, hiring discrimination, racial inequality, labor market frictions, unemployment, wage gaps, wealth distribution, business cycles
@@ 1 Introduction
Market competition and increased access to information are expected to eliminate discrimination over time. Nonetheless, substantial empirical evidence shows that discriminatory practices persist across markets, perpetuating significant racial disparities. [^1] For example, Couch and Fairlie (2010); Biddle and Hamermesh (2013); Kuhn et al. (2020); Derenoncourt and Montialoux (2021); Lippens et al. (2023); Lang and Spitzer (2020). Resume studies consistently reveal racial discrimination even when controlling for candidate qualifications (e.g., Bertrand and Mullainathan, 2004; Bertrand and Duflo, 2017; Quillian et al., 2017; Kline et al., 2022). Gaps in labor income and wealth between black and white households have persisted long after the Civil Rights Movement (e.g., Cajner et al., 2017; Derenoncourt et al., 2023). Small and Pager (2020) argues that discrimination reinforces itself across domains and constitutes a form of market failure. Only recently has macroeconomic research begun exploring racial disparities. [^2] See, for example, Nakajima (2021), Aliprantis et al. (2023), Boerma and Karabarbounis (2021), Ganong et al. (2020), Lee et al. (2021). Yet, the role of discriminatory practices in shaping these disparities remains underexplored. Developing a general equilibrium framework to explain the persistence of discrimination and its consequences is crucial for understanding the macroeconomic impacts of racial disparities and informing effective policymaking.
This paper constructs a search-and-matching model to examine the macroeconomic effects of persistent racial discrimination in hiring. We find that hiring discrimination explains roughly 57% of the black-white unemployment rate gap, 48% of the average wage gap, and 16% of the median wealth gap in the steady state. Moreover, discriminatory hiring amplifies the volatility of unemployment and wages among black workers over the business cycle. Eliminating hiring discrimination not only improves the welfare of black workers but also enhances overall economic welfare.
@@
The model features two types of firms: prejudiced firms, which hire only white workers, and non-prejudiced firms, which hire all workers equally. As a result, all workers compete for jobs at non-prejudiced firms, but white workers have the additional opportunity to match with prejudiced firms. Matched worker-firm pairs negotiate wages to maximize their joint surplus. Upon calibration, the model endogenously produces a lower job-finding rate, higher unemployment rate, and lower bargained wages for black workers. Even without additional financial frictions, the model generates substantial racial disparities in wealth accumulation.
@@
The main message of this paper is that labor market frictions can sustain hiring discrimination as an equilibrium outcome. Our calibration shows that prejudiced firms incur higher vacancy posting costs to discriminate against black workers. Yet, they sustain positive profits by maintaining a lower turnover rate upon hiring white workers. Moreover, hiring discrimination affects the equilibrium economy through firms’ vacancy postings. Prejudiced firms compete with non-prejudiced firms for white workers, driving up their wage rates across the economy. This wage pressure reduces the expected profit of non-prejudiced firms, leading them to post fewer vacancies overall. The resulting decline in job openings at non-prejudiced firms directly limits employment opportunities for black workers, contributing to lower wages, higher unemployment, and adverse wealth accumulation. In a counterfactual economy without prejudiced firms, these racial disparities narrow substantially.
Eliminating discriminatory hiring also yields significant welfare gains. Black workers experience a 12.9% increase in welfare as non-prejudiced firms expand their vacancy postings, which in turn improves their job opportunities and wages. Although white workers incur a modest 0.5% welfare decline due to the loss of exclusive job opportunities, the overall welfare of the economy rises by 1.5% because the gains for black workers more than offset the losses for white workers.
We further incorporate aggregate shocks into our benchmark economy to examine how hiring discrimination shapes business cycle dynamics. Our model replicates key empirical patterns: a countercyclical and volatile racial unemployment gap, as well as a more procyclical and volatile wage response among black workers. In a counterfactual model without prejudiced firms, the volatility of unemployment rates and wages for black workers is substantially reduced relative to white workers. Moreover, hiring discrimination amplifies the volatility of racial disparities in consumption and wealth over the business cycle. These findings suggest that firm-side hiring discrimination plays a crucial role in driving the distinct business cycle dynamics of black and white workers, which cannot be fully explained by demographic factors alone (Cajner et al., 2017; Boulware and Kuttner, 2019, 2024).
@@
This paper contributes to the ongoing discussion on the persistence of racial inequality by focusing on the disparate labor market conditions of black and white workers. Numerous studies have documented racial discrimination in pay and employment opportunities that persist despite policy interventions (e.g., Coate and Loury, 1993; Black, 1995; Rosén, 1997; Bertrand and Mullainathan, 2004; Manduca, 2018). Recent empirical research stresses that racial wage and employment gaps persist across education levels, skill sets, and cohorts. These disparities are not solely attributable to differences in individual qualifications (e.g., Pena, 2018; Cheng et al., 2019). Building on this insight, our paper develops a model in which discriminatory hiring practices endogenously perpetuate labor market inequalities. Our work adds to the literature by providing a theoretical framework that explains how hiring discrimination can be sustained as an equilibrium outcome in the labor market.
@@
While much of the literature focuses on static disparities, fewer studies document how macroeconomic fluctuations impact black and white workers differently. Notably, Couch and Fairlie (2010) demonstrate that black workers are disproportionately affected during recessions, being the last hired in upturns and the first fired in downturns. Biddle and Hamermesh (2013) find that the discriminatory wage gap is procyclical, widening during economic expansions. Similarly, Cajner et al. (2017) show that black workers experience higher unemployment rate volatility and are more likely to be involuntarily employed part-time. Daly et al. (2020) suggest that limited employment opportunities for black workers contribute to the growing racial earnings gap. Recent research by Boulware and Kuttner (2019, 2024) further highlights that racial disparities intensify during downturns, driven by increased hiring discrimination and differential sensitivity of black employment to cyclical fluctuations, effects that cannot be fully attributed to black workers’ concentration in cyclically sensitive industries. Our paper adds to this body of work by evaluating the impact of hiring discrimination on the cyclical aspects of racial disparities, examining how discriminatory practices exacerbate the volatility and cyclicality of employment and wages for black workers.
@@
An emerging strand of literature explores racial disparities in wealth accumulation. Studies such as Derenoncourt et al. (2023), Derenoncourt et al. (2022), Kuhn et al. (2020), Barsky et al. (2002), and McIntosh et al. (2020) document significant wealth gaps between black and white households. For example, Derenoncourt et al. (2023) provides a historical account of wealth segregation over the past 150 years, showing persistent and substantial disparities. Boerma and Karabarbounis (2021) and Aliprantis et al. (2023) analyze the impact of historical discrimination on earnings, bequests, and capital returns within steady-state models that do not consider aggregate risks. Furthermore, Ganong et al. (2020) demonstrates that income risks affect individuals differently across racial groups due to wealth differences. Building on these findings, Bartscher et al. (2021) and Lee et al. (2021) discuss how monetary policy can have disparate effects on workers of different racial backgrounds. Our paper contributes to this literature by showing that employer discrimination not only affects immediate labor market outcomes but also spills over to long-term wealth accumulation disparities between black and white workers.
@@
Germane to our project, Nakajima (2021) develops a search-and-matching model to examine how monetary policy exacerbates racial differences in the labor market. Our model explicitly distinguishes between prejudiced and non-prejudiced firms in the hiring process and directly examines how hiring discrimination impacts black workers and the broader economy. We analyze the spillover effects of discriminatory hiring practices on both labor and wealth disparities, offering a comprehensive theoretical understanding of these dynamics over the long run and across business cycles.
@@
The rest of the paper proceeds as follows. Section 2 lays out the theoretical framework. Section 3 discusses the calibration strategy. Section 4 examines the steady-state implications of racial discrimination. Section 5 discusses the business cycle implications. Section 6 concludes the paper.
@@ 2.1 Environment
The model has a unit measure of workers, who are either black or white, $R\in\{bl,wh\}$. Two types of firms post vacancies subject to search and matching frictions. Prejudiced firms ($p$) discriminate in hiring by effectively excluding black applicants, whereas non-prejudiced firms ($np$) hire workers without regard to race. A matched firm-worker pair corresponds to a filled job position, making firms and jobs synonymous within our model. Black workers thus receive offers exclusively from non-prejudiced firms, while white workers can match with both types of firms, increasing their employment opportunities. This assumption closely aligns with extensive empirical evidence on racial disparities in hiring. Audit studies by Bertrand and Mullainathan (2004) find that resumes with white-sounding names receive approximately 50% more callbacks than identical resumes with black-sounding names. Similarly, a meta-analysis by Quillian et al. (2017) confirms that white applicants consistently receive, on average, 36% more callbacks. Although discrimination occurs in other aspects such as wages, promotions, and separations, we specifically focus on hiring discrimination due to its documented significance, allowing us to clearly evaluate its distinct macroeconomic effects.
Workers are either employed (with $p$ or $np$ firm) or unemployed. Those who become unemployed in the current period receive unemployment benefits and continue to receive them in the future with probability $P_{e}$. The workers face idiosyncratic productivity shocks $s$, following an AR(1) process $\log(s^{\prime})=\rho_{s}\log(s)+\epsilon_{s}$, with $\epsilon_{s}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{s}^{2})$. [^3] To ease exposition, we drop the time subscripts and use a prime symbol (’) to denote variables in the next period. Workers also differ in their asset holdings. They can save using risk-free assets to partially insure themselves against the labor market risks. Following Mukoyama (2013), workers also receive race-specific extreme wealth shocks with probability $\epsilon_{R}\in\{\epsilon_{bl},\epsilon_{wh}\}$. Upon realization of the shock, a worker loses all their wealth. Altogether, workers are heterogeneous across race ($R$), labor market status ($e$), idiosyncratic productivity ($s$), and wealth ($a$). The endogenous distribution of workers is summarized by $\mu(e,R,s,a)$. In the steady state, individual workers may move around the distribution, but the overall distribution remains stationary. Thus, $\mu$ is not a state variable in the steady state model but becomes an aggregate state variable in the augmented model with aggregate shocks discussed in Section 5.
@@ 2.2 Labor market search and matching
The total number of unemployed workers denoted as $u$, is the sum of unemployed black workers $(u_{bl})$ and unemployed white workers $(u_{wh})$. There are $v_{np}$ vacancies available in the non-prejudiced market, while the number of vacancies in the prejudiced market is $v_{p}$. Non-prejudiced firms search among both black and white unemployed workers, resulting in a non-prejudiced market tightness defined as $\theta_{np}=\frac{v_{np}}{u}$. In contrast, prejudiced firms only hire unemployed white workers, which gives the prejudiced market tightness $\theta_{p}=\frac{v_{p}}{u_{wh}}$. Following the works of Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018), unemployed workers and vacant firms match through a constant returns to scale matching function
$M(u,v)=\frac{uv}{(u^{\iota}+v^{\iota})^{1/\iota}}$ (1)
with $\iota>0$. As documented by Den Haan et al. (2000), this functional form ensures that the matching probabilities stay within 0 and 1. The probability for an unemployed worker to match with a vacant $np$ firm is $f(\theta_{np})=M(u,v_{np})/u=(1+\theta_{np}^{-\iota})^{-1/\iota}$, while the probability that a white unemployed worker matches with a vacant $p$ firm is $f(\theta_{p})=M(u_{wh},v_{p})/u_{wh}=(1+\theta_{p}^{-\iota})^{-1/\iota}$. Correspondingly, the probability of filling a vacant $np$ firm is $q(\theta_{np})=M(u,v_{np})/v_{np}=(1+\theta_{np}^{\iota})^{-1/\iota}$, while the probability of filling a vacant $p$ firm is $q(\theta_{p})=M(u_{wh},v_{p})/v_{p}=(1+\theta_{p}^{\iota})^{-1/\iota}$. In addition, non-prejudiced matches get destructed with probability $\lambda_{np}$, while the prejudiced matches separate at the rate $\lambda_{p}$.